Peeblesshire Beekeepers Association

Association Apiary

PBKA Pub Meeting 6th December

The last meeting of the PBKA this year will be on Wednesday 6th Dec at 7.30pm in the back bar of the Neidpath Inn in Peebles.

If you are thinking of starting beekeeping in the new year we can advise on beginners classes and meetings that will help.

All welcome.

PBKA Pub Meeting 8th November

Only a few apiary jobs to do at this time of year – making hives weather and mouse proof, checking stores and planning varroa treatment.

If you want advice about planning for next season, or to get some exhibits ready for this month’s honey show, or if you are a novice and just want to find out more, then come along and chat to the local beekeepers.

This months meeting will be on Wednesday 8th Nov at 7.30pm in the back bar of the Neidpath Inn in Peebles.

PBKA Pub Meeting 4th Oct

It’s the time of year to think of winter preparations for the bees.

Maybe its time to be extracting and preparing honey if you’ve been lucky with a honey crop this year.

Also a chance to review the season and think about next year.

As usual all welcome to discuss this and all things beekeeping on Wednesday 4th Oct at 7.30pm in the back bar of the Neidpath Inn in Peebles.

See you there.

Apiary Visit 10th September Report

Richard L and Peter went to the apiary on Sunday 10th September.

The weather hadn’t looked promising but it was actually fine and even sunny while we there.

The guys doing the walling were also there and had reached the section where the storage huts are. They needed to get their scaffolding in at the wall. So with their help, we shifted the huts to behind the line of hives and repacked the contents. Very kindly they offered to do the bulk of the heavy lifting and we didn’t argue too much!

There are only 2 hives in the central area with bees now and we topped up the feeders on both. One had a super on which we removed along with the used up MAQ’s strips. This one had quite a large varroa count which is hopefully still mites falling from the treatment. All the hives need to be checked for varroa now and judged to see what treatment they are going to get. The third active hive was in at the wall. Since the wallers were actively working here and had the compressor going adjacent to the hive we didn’t open this one up. Peter may go back and check this one if there is a warm evening this week.

The walling work may be completed this week and the guys have just worked around the bees, so in the end there wasn’t any need to relocate the hives. Longer term, it would be nice to find out more about the plans for the walled garden but at the moment this information is hard to come by.

News from the out apiary at Merlindale – Peter went out to Dankas on Saturday. The 2 association hives are doing very well. One which had no laying queen has brood again now and plenty bees. There are also good stores and we left the super on that one. The other hive which has still a huge amount of bees had also filled the best part of 2 supers. We took one of these off for association honey and left the remaining one for the bees. We moved Dankas colony from the association hive into her own poly hive and were amazed at how much brood and bees there were. All were good tempered with us although they had disgraced themselves earlier by stinging the gardener who had been strimming in front of the hives. He wasn’t too put out fortunately and will take some more precautions the next time.

Finally thanks to Brian for the photos taken earlier in the season and I’ll post some more of these in due course.

At work in the apiary.

A shallow frame has been in the brood box.

Apiary Visit 10th September

Apiary meeting this Sunday 2pm, even if the weather is poor – Peter will be there for a short time only, to check the colonies, re-organise hive 1 (if weather isn’t too bad), feed hive 2, and possibly do some anti-varroa treatment.

If you need to borrow a beesuit, please contact Peter.

There will be no apiary visit on 17th.

Information about the winter programme will follow shortly.

Apiary Visit 20th August Report

Apiary visit 20th August.

Must be holiday time! It was a relief to get some warm dry weather at last, and a chance to go through the hives. Mandy, Claire and Shirley checked out the reduced number we have now (mostly due to relocation, and nucs going out to this year’s beginners), every colony different, so it turned out to be a useful teaching session.

The small nuc near the wall was coming on well, had filled 4 frames, so we put it into a proper hive. An empty hive was placed just in front of it, then Claire and Shirley transferred the frames. The bees were somewhat confused to start with, then settled down. They didn’t seem short of stores (given their size – they’d been fed generously until a fortnight ago) and with the weather forecast to improve, they were safe to forage for themselves without extra feeding.

Hive 1 had created a new queen in late June, after the queen and some frames of bees were removed to create a nuc. We hadn’t seen any brood on the last 2 inspections, and this time no brood again, but lots of drones and scattered drone brood, so we suspected a drone-laying worker (as some were above the queen excluder). In this situation, the colony is effectively a write-off without a queen, but the remaining bees can be used to boost another colony. A frame of eggs and brood was inserted from hive 3, to boost the worker numbers, and attract any remaining workers to a single frame, then the nuc was opened up again, and a “paper marriage” carried out. Paper was laid across the frames in the nuc’s brood box, holes made in the paper, and hive 1 was placed on top. The scent from the 2 colonies should slowly mingle, and the 2 lots of bees should combine without fighting.

This colony (next to the storage unit) will need re-assembling within a week, and fed syrup until the remaining combs in the brood box have been drawn out.

Hive 2 was in a sorry state – no brood was seen in the previous inspection (it needed to make a new queen after frames were removed to create a nuc): this week a full super of honey had gone, still no brood visible, but no drone brood either, so probably the queen is still here. This been a difficult summer, with some colonies running out of food. Hopefully this is what’s happened here – the queen is still present, but not laying due to lack of food. Easily remedied, we hope, with lots of feeding. Syrup was put on the hive – this will need checking within a week, and more syrup added as needed.

Hive 3: had had a very heavy varroa infestation: this had been treated about 3 weeks ago (MAQ strips) with spectacular results – heaviest mite fall I’ve ever seen. This week – with the MAQ strips still present – there was still a mite drop, but probably only a Daily Mite Drop (DMD) of about 4. The strips were left in place. The colony only had 3 frames of stores in the super, so a feeder and syrup was added, to let them get started and help boost the last of this summer’s foraging. The queen was laying well, with plenty of brood, larvae and eggs seen.

The last colony to check was a swarm collected this summer: it was very strong at the last inspection, and this time it had obviously outgrown its five-frame nuc box, so was swiftly transferred to a proper hive. A feeder and syrup was added to help them drawn out the additional 6 frames they now had. As there was only a small amount of syrup left, this colony (closest to the trees) will need feeding again before the weekend.

This visit – with so much variability between only 5 colonies – serves as a reminder not to take anything for granted, to check colonies for stores, look out for robbing by wasps of other bees, check the mite drop and work out the end-of-season varroa treatments.

The heather probably has only a week or so left (it opened 7-10 days early this year), so there will be little forage about except for rosebay willowherb, possibly dandelions if they get a chance of a second flowering, and Himalayan Balsam. The latter is an invasive alien, so probably its only fans are bees who relish the copious nectar, but don’t seem to like the white pollen. When they’ve been working the Himalayan Balsam they come back powdered with white pollen (“ghost bees”) that they don’t bother gathering onto their pollen baskets.

Winter bees: colonies are starting to raise the bees that will take them right through the winter, instead of having the 6-week lifespan of the summer bees. Make sure that all colonies have plenty of food at this stage, to secure the colony strength through the winter and into spring.

This year’s honey crop looks likely to be very poor: it will be interesting to see how much heather honey comes in.

Apairy Visit 30th July

Sunday 30th July: apiary visit – although the forecast doesn’t look good so far. If weather permits, we’ll be at the Association apiary 2-4; if we have to cancel I’ll email at 11am, and will try to arrange a midweek visit instead.
As usual, please wear clean beesuits, gloves etc (let me know in advance if you need to borrow one), and perhaps bring some refreshments to share.

Mandy

Apiray Visit 16th July 2017

Summary of apiary visit 16th July 2017

Apiary 16th July

Hive 1: 16th July was the earliest day we could expect the new queen to lay, but there were no eggs or larvae yet – check again next week. A large number of drones were present; bees seemed calm.

Next time: check for eggs/larvae – if new queen seen, mark her. Check for stores, pollen being brought in. Consider adding a frame of eggs if no sign of a laying queen.

Hive 2: the queen had been removed on 6th June to create a nuc: stores were OK, no eggs or larvae seen yet, bees seemed calm. Check again next week.

Next time: check for eggs/larvae, a laying queen. Consider adding a frame of eggs if no sign of a laying queen.

Hive 3: a high mite drop again; but excellent brood pattern, this hive would be a good source of eggs for hives 1 or 2, if need be – hopefully it won’t be necessary, as that would then mean 3 of our hives were from the same genetic line, which could be a bit risky. Syrup added as stores were running low, and it has a large population to feed.

Next time: Peter Stevenson volunteered to treat the hive with MAQ during the week. Check mitre drop.

Hive 6: at a new location – checked by Peter and David

Doing well – brood at all stages on 7-8 frames. Plenty stores. Bees calm. Primed queen cells noted. Needs to be inspected again soon.

Hive 7: at a new location – checked by Peter and David

No brood apart from 2 sealed and  possibly unsealed queen cells noted. 1 sealed cell knocked down to reveal a developing queen. Very puzzling!

2 frames of brood with eggs taken from Hive6 and given to Hive7 on either side of the Queen cell – all marked by pins. Intention to slow down development of hive6 to discourage swarming and to check for queenlessness in Hive 7.

At least one nuc yet to be created from these hives: reserved.

(The empty hive 7 is with AC for cleaning, along with 3 empty nuc boxes)

Swarm 1 was bought and taken away by Duncan on Monday

Swarm 2: put in brood box, and syrup feed added.

Next time: check if ready to go.

Nuc 2: feed added, quite a small colony, only on 2 frames, but sealed worker brood seen, it should continue to build up. Continue to monitor and feed as necessary.

Nuc 3: bought and taken away by Ilknur and Brian.

Kailzie colonies: all doing well, 2 nucs available, a further nuc reserved.

 

Colonsay native honeybee colonies have been relocated; both have developed chalkbrood, possibly an indication that the previous location had insufficient forage and was too damp.

They seem to be thriving at their new site.

A further nuc at AC’s apiary, ready to go.

One of the nucs has been reserved by Em and Kelly, the other by Fiona MacAuley: this leaves potentially 3 nucs available.

For the next apiary visit: more record sheets needed; syrup needed; feeders have been taken away for cleaning and need to be returned. Syrup needed for small nuc.

There have been many reports from local beekeepers about problems of getting new queens established, and of queenless colonies rejecting both frames of eggs, new queens and even queen cells. For some reason it seems to be taking longer for queens to start laying.

PBKA Pub Meeting 5th July 2017

July is often when a local honey crop can be had in the Peebles area. Lime trees will be flowering this month and hopefully providing the bees with good forage.

Keep up to date with the local beekeepers and find out how the season is going.

Some Association nucs have gone to new homes and other colonies may need temporary or permanent homes.

The native bee project is under way.

Lots to discuss on Wednesday 5th July at 7.30pm in the back bar of the Neidpath Inn in Peebles.

 

Apiary Visit 25th June 2017

2017 06 25 Apiary visit

Only 3 this time: Mandy, Julia and Shirley. Anne had hoped to come, but a puncture en route meant she didn’t make it. Still, a very useful session, lost for the beginners to see, they both handled the bees and had plenty of questions to ask. And many thanks to Shirley for the lovely cake!

Hive 1: no brood, a large number of drones and many queen cells – most open, a couple torn down, and one resealed. Bees were quiet but seemed happy, so assumed they had a queen that was not yet laying: the super beneath the brood box was removed, all the bees brushed gently into the brood box, to which (just in case) a frame of eggs was added from the large well-established swarm. Two old frames were replaced with fresh foundation, then the supers were replaced – one was full, one had drawn comb, and another super was optimistically added on top. (Well, the lime should be starting mid-July – they have enough stores to last till then.)

Next time: no further disturbance for 3 weeks please! Daily mite drop: 0.5

Hive 2: The brood box was not disturbed, as waiting for a new queen to start laying. It had a brood box above the excluder in the hope that this would be drawn out as a supply of drawn comb, but the bees (after 4 weeks) had declined to do this, so it was removed. The super above it was almost full, so another super was added. DMD: 0.8

Next time: check for brood/eggs. Mark queen if seen.

Hive 3: a frame of irregular comb was moved to the edge of the brood box – this should be removed and replaced as soon as the brood has hatched. A queen excluder was added.

DMD: 3.2

Next time: replace malformed brood frame if possible.

Hive 5 and 6 are away from home at the moment – last inspected 17th June, so looking for volunteers to go out and check them up.

Swarms and nucs:

Swarm 1 on 9 frames (one removed to go into Hive 1): this swarm will go to a new home shortly, prepare for moving.

Swarm 2: needs to go in a brood box next time. Check syrup, add queen excluder and feed.

Sterilise a new brood box and fill with frames, ready for use.

Nucs 2 and 3 – both doing well, but still only on 2 or 3 frames.

Check syrup, replace feeders on both to keep mould-free.